Question 3: Headsound manufactures headphone cases. Question 3: HeadSound manufactures headphone cases. During September 2014, the company produced and sold 107,000 cases and re
Answer #1
1)
D.mat price variance= AQ*SP-AQ*AP
D.mat price variance= 209000*0.15-209000*0.20
D.mat price variance= $                                                  (10,450.00) UnFavourable
D.mat usage variance= SQ*SP-AQ*SP
D.mat usage variance= (2.00*107000)*0.15-(209000*0.15)
D.mat usage variance= $                                                         750.00 Favourable
Total Direct Materials Cost Variance= SQ*SP-AQ*AP
Total Direct Materials Cost Variance= (2.00*107000)*0.15-(209000*0.20)
Total Direct Materials Cost Variance= $                                                    (9,700.00) UnFavourable
D.LABOR RATE variance= AH*SR-AH*AR
D.LABOR RATE variance= 1640*8-1640*8.15
D.LABOR RATE variance= $                                                       (246.00) UnFavourable
D.LABOR usage variance= SH*SR-AH*SR
D.LABOR usage variance= (107000*0.02)*8-1640*8
D.LABOR usage variance= $                                                      4,000.00 Favourable
Total Direct LABOR Cost Variance= SH*SR-AH*AR
Total Direct LABOR Cost Variance= (107000*0.02)*8-1640*8.15
Total Direct LABOR Cost Variance= $                                                      3,754.00 Favourable
2)
Variable OH rate variance= AH*SR-AH*AR
Variable OH rate variance= 1640*11-8000
Variable OH rate variance= $                                                   10,040.00 Favourable
Variable OH effeciency variance= SH*SR-AH*SR
Variable OH effeciency variance= 107000*0.02*11-1640*11
Variable OH effeciency variance= $                                                      5,500.00 Favourable
Total Variable OH variance= SH*SR-AH*AR
Total Variable OH variance= 107000*0.02*11-8000
Total Variable OH variance= $                                                   15,540.00 Favourable
Fixed cost volume variance= Budgeted OH-actual units*budgeted rate
Fixed cost volume variance= 31040-107000*0.32
Fixed cost volume variance= $                                                    (3,200.00) Favourable
Budgeted rate= 16*0.02
Budgeted rate= $                                                             0.32
Fixed cost variance= Actual OH-absorbed OH
Fixed cost variance= 30000-107000*0.32
Fixed cost variance= $                                                    (4,240.00) Favourable

3) As company has used better quality materials it should result in less usage of materials but prices would be higher of better quality materials so

Effect on

Material price variance would be Unfavourable

Material usage variance would be Favourable

Please like the solution if satisfied with the answer and if any queries please mention it in comments...thanks